The combinations on the lockers at the high school consist of three numbers, each ranging from 0-39. If a combination is chosen at random, what is the probability that the first two numbers are multiples of 9 and the third number is a multiple of 4?
Answer #1
There are four multiples of 9 and 9 multiples of four between 0 and 39.

If each number is independently chosen from 0 to 39, the probability of what you describe is
4/39 x 4/39 x 9/39 = 144/55,319 = 0.002428

or about 1 in 412
answered by: Jay Tillman
Answer #2

7/1280

Correct answered by Mathlab